OpenStudy (anonymous):

Help is appreciated(: How would I solve 2cosx+5=4?

OpenStudy (anonymous):

you want to solve for cos x?

OpenStudy (anonymous):

I believe so.

OpenStudy (anonymous):

Subtract 5 from both sides, you get 2 cos x = -1 Now divide both sides by 2, you get cos x = - 1/2

OpenStudy (anonymous):

thanks! so for 1/2-sinx=1, would I have to subtract the 1/2? @Easyaspi314

OpenStudy (anonymous):

Yes, but then you get - sin x = 1/2 Then you have to divide both sides by -1, so you get sin x = - 1/2

OpenStudy (anonymous):

Okay, thanks again(:

OpenStudy (anonymous):

sure.